OpenStudy (anonymous):

whats the solution of the proportion 6/10=9X

OpenStudy (anonymous):



OpenStudy (anonymous):

YOUR ANSwer is \[x = 6/9\]

OpenStudy (anonymous):

A. 15 B. 5.4 C. 30 D. 13 it has to be one of these doe.

OpenStudy (anonymous):

did you type in the equation right?

OpenStudy (anonymous):

6 over 10 equeals 9 over X

OpenStudy (anonymous):

this is all the way wrong... at the first answer 6/10 = 9x 3/5 = 9x 9x = 3/5 9x/9 = 3/5/9 x = 1/15 so 15 would be your answer